OpenStudy (anonymous):

Find the point(s) of intersection of the parabola with equation y = x2 - 5x + 4 and the line with equation y = 2x - 2

OpenStudy (sohailiftikhar):

it means you have to find the slution set

OpenStudy (campbell_st):

all you need to do is equate them so \[2x -2 = x^2 -5x + 4\] then you can rewrite the the equation as \[x^2 - 7x + 6 = 0\] this can now be solved by factoring... to get the x values of each point subsitute the x values into either equation to find the corresponding y values hope it helps

OpenStudy (anonymous):

thank you @campbell_st